# Spin Chern number in altermagnets

### v1

@@ III.2 $\mathbb{Z}_{2}$ bulk invariant for $C_{4}\mathbb{T}$ on 2D torus
$r\otimes\mathbb{Q}:\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})\otimes\mathbb{Q}\stackrel{{\scriptstyle\cong}}{{\to}}\left(K_{C_{2}}^{0}(T^{2})\otimes\mathbb{Q}\right)^{\mathbb{Z}_{2}}$ (46)
implies in particular that any integer invariant of $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})$ could be detected with the integer invariant in $K_{C_{2}}^{0}(T^{2})^{\mathbb{Z}_{2}}$ given by the restriction map $r$.
The bundles that generate the bulk invariants of $K_{C_{2}}^{*}(T^{2})$ are well known and can be constructed. Let $\mathcal{H}$ be the line bundle over $T^{2}$ with $C_{2}$ action whose Chern number is $1$. We could take as $\mathcal{H}$ the line bundle given by the valence band of the Hamiltonian
@@
with $C_{2}$ action $(k_{x},k_{y})\mapsto(-k_{x},-k_{y})$ and associated matrix $\mathrm{diag}(i,-i)$. Here $C_{2}$ has for eigenvalue $i$ on $\Gamma$ and $-i$ on the other three TRIMS. The Chern number $c_{1}(\mathcal{H})$ of $\mathcal{H}$ is $1$. Its conjugate $\widehat{\mathcal{H}}$ has the opposite action of $C_{2}$, its Chern number $c_{1}(\widehat{\mathcal{H}})$ is $-1$ and the $C_{2}$ eigenvalues are the complex conjugate of the ones of $\mathcal{H}$.
The tensor power $\mathcal{H}^{\otimes 2n+1}$ of $\mathcal{H}$ is also a $C_{2}$-equivariant line bundle and the Chern number distinguishes the line bundles $\mathcal{H}^{\otimes 2n+1}$ in $K_{C_{2}}^{0}(T^{2})$.
@@
$\mathcal{H}^{\otimes 2n+1}\oplus\widehat{\mathcal{H}}^{\otimes 2n+1}\in K_{C_{2}}^{0}(T^{2})^{\mathbb{Z}_{2}}$ (48)
become all equal in $K_{C_{2}}^{0}(T^{2})$. Note that the Chern number of the sum $\mathcal{H}^{\otimes 2n+1}\oplus\widehat{\mathcal{H}}^{\otimes 2n+1}$ is zero, and they all share the same eigenvalues on the TRIMs. In other words, the bulk invariant of the bundles $\mathcal{H}^{\otimes 2n+1}\oplus\widehat{\mathcal{H}}^{\otimes 2n+1}$ is zero, and because of this, the Chern number for each separate piece cannot be extracted from the K-theory group.
@@
$\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2},$ (49)
where the $\mathbb{Z}_{2}$ invariant is precisely the bulk invariant that is of interest. This bulk invariant is precisely the Chern number (mod $2$) of one of the pieces of its decomposition in the generators of $K_{C_{2}}^{0}(T^{2})^{\mathbb{Z}_{2}}$. But since we cannot extract Chern numbers from $K_{C_{2}}^{0}(T^{2})^{\mathbb{Z}_{2}}$, we are going to add the symmetry of the spin $z$ operator in order to extract this Chern number.
Let $S_{z}$ be the spin $z$ and let us assume that the spin acts on the bundles defining $\mathcal{K}^{*}_{C_{4}\mathbb{T}}(T^{2})$. This happens for example when the Hamiltonian commutes with the spin. Incorporating $S_{z}$ to the group of symmetries we have that $\{S_{z},C_{4}\mathbb{T}\}=0$ and that $[S_{z},C_{2}]=0$. Now let us consider the following commutative diagram of restrictions:
where the left vertical maps forget the $S_{z}$ action, the left horizontal maps are the restriction to the complex part defined in Eqn. (39), and the right vertical map is the Chern number of the spin-up part. Here $\mathcal{K}^{0}_{C_{4}\mathbb{T},S_{z}}(T^{2})$ and $K^{0}_{C_{2},S_{z}}(T^{2})$ denote the equivariant magnetic and complex K-theories associated to the groups $\langle C_{4}\mathbb{T},S_{z}\rangle$ and $\langle C_{2},S_{z}\rangle$ respectively.
@@
Denote by $E$ the vector bundle that the valence bands of the altermagnetic Hamiltonian of Eqn. (1) defines. We know that for $G^{2}<M_{0}<2$ (say $M_{0}=1.5$), its restriction to the K-theory of $C_{2}$ splits as the line bundles defined by the valence bands for each Hamiltonian block; in other words
$r(E)=\mathcal{H}\oplus\widehat{\mathcal{H}}\in K_{C_{2}}^{0}(T^{2})^{\mathbb{Z}_{2}}.$ (55)
@@
Now we can finish our argument. In order to detect the Chern number of one of the pieces of the class $\mathcal{H}\oplus\widehat{\mathcal{H}}$ in the K-theory ${K}_{C_{2}}^{0}(T^{2})$ we use the spin $z$ splitting.
Remark: Note that we cannot extract the Chern number of one of the pieces in $\mathcal{H}\oplus\widehat{\mathcal{H}}$ from its image in ${K}_{C_{2}}^{0}(T^{2})$. In ${K}_{C_{2}}^{0}(T^{2})$ the element $\mathcal{H}\oplus\widehat{\mathcal{H}}$ has no bulk invariant.
Incorporating the spin, we see that the vector bundles in $K^{0}_{C_{2},S_{z}}(T^{2})$ can be split into spin-up and spin-down thus having:
$K^{0}_{C_{2},S_{z}}(T^{2})\cong K^{0}_{C_{2}}(T^{2})^{\uparrow}\oplus K^{0}_{C_{2}}(T^{2})^{\downarrow}.$ (61)
Since the $\mathbb{Z}_{2}$ action on vector bundles on $K^{0}_{C_{2},S_{z}}(T^{2})$ flips the spin (this follows from the fact that $C_{4}\mathbb{T}$ anticommutes with $S_{z}$), the $\mathbb{Z}_{2}$-invariant part can be parametrized with the spin-up bundle:
$K^{0}_{C_{2},S_{z}}(T^{2})^{\mathbb{Z}_{2}}\cong K^{0}_{C_{2}}(T^{2})^{\uparrow}.$ (62)
@@
This splitting allow us to extract only the component of the spin-up, and from this component we may take its Chern number; this is the spin-up Chern number. Knowing that this spin-up Chern number is odd, permits us to distinguish the bulk invariant of the bundle in $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})$.
From diagrams (54) and (60) we can conclude the following. The bundle $E$ generates an integer invariant in $\mathcal{K}^{0}_{C_{4}\mathbb{T},S_{z}}(T^{2})$ since it maps to an integer invariant in ${K}^{0}_{C_{2},S_{z}}(T^{2})^{\mathbb{Z}_{2}}$. This follows from the fact that the restriction map in the upper left of diagram (54) induces an isomorphims rationally.
We now claim that the bundle $E$ generates the $\mathbb{Z}_{2}$-bulk invariant in $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})$. For this we need to determine explicitly the complete K-theory invariants in ${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{2})$. This will be done in the next section.
@@ III.2.1 Subdivision of the torus with 4 squares
To find the magnetic equivariant K-theory invariants in ${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{2})$ we will make use of an equivariant subdivision of the torus. The long exact sequence of the pairs will allows to find the desired K-theory groups.
@@
$\mathcal{K}^{even}_{C_{4}\mathbb{T}}(X_{0})\cong$ $\mathbb{Z}\langle\bar{\Gamma},\bar{M},\bar{X}_{+},\bar{X}_{-}\rangle$ (66)
where both $\bar{\Gamma}$ and $\bar{M}$ denote the 2-dimensional representations of $\langle C_{4}\mathbb{T}\rangle$ on $\Gamma$ and $M$ respectively and $\bar{X}_{+}$ and $\bar{X}_{-}$ denote the 1-dimensional representations of $C_{2}$ in $X$ (recall that $(C_{2})^{2}=-1$).
@@
$\mathcal{K}^{odd}_{C_{4}\mathbb{T}}(X_{1},X_{0})\cong$ $\mathbb{Z}\langle\bar{e},\bar{h}\rangle$ (67)
$\mathcal{K}^{even}_{C_{4}\mathbb{T}}(X_{1},X_{0})\cong$ 00 0 (68)
$\mathcal{K}^{even}_{C_{4}\mathbb{T}}(X_{2},X_{1})\cong$ $\mathbb{Z}\langle\bar{Z}\rangle$ (69)
$\mathcal{K}^{odd}_{C_{4}\mathbb{T}}(X_{2},X_{1})\cong$ 00 0 (70)
which follows from the fact that the $C_{4}\mathbb{T}$ action on the 1 and 2-cells is free.
@@
$0\to\mathcal{K}^{-2}_{C_{4}\mathbb{T}}(X_{1})\to\mathcal{K}^{-2}_{C_{4}\mathbb{T}}(X_{0})\stackrel{{\scriptstyle\delta_{0}}}{{\to}}\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1},X_{0})$ (71)
$\to\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1})$ $\to 0.$
@@
$\mathcal{K}^{-2}_{C_{4}\mathbb{T}}(X_{0})\stackrel{{\scriptstyle\delta_{0}}}{{\to}}$ $\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1},X_{0})$ (72)
@@
$\mathcal{K}^{-2}_{C_{4}\mathbb{T}}(X_{0})=\widetilde{\mathcal{K}}_{C_{4}\mathbb{T}}(\Sigma^{2}X_{0})$ (75)
and therefore we can imagine the generators $\bar{\Gamma},\bar{M},\bar{X}_{+},\bar{X}_{-}$ to be virtual bundles over $S^{2}$. Since the conjugation action on $K^{0}(S^{2})$ maps the Hopf bundle to its inverse, we have that the restriction map $\mathcal{K}^{-2}_{C_{4}\mathbb{T}}(\{\Gamma,M\})\to{K}^{-2}(\{\Gamma,M\})$ is trivial. The image of $\bar{X}_{+}$ and $\bar{X}_{-}$ depends on the orientation of the 1-cells that attach to the point $X$. The 1-cell $h$ leaves from $X$ and the 1-cell $e^{\prime}$ arrives, but we need to use the automorphism that maps the information on $e^{\prime}$ to $e$ by conjugation, that on this dimension is given by multiplication by $-1$ (again the Hopf bundle is mapped to its inverse). Therefore we have
$\mathcal{K}^{-2}_{C_{4}\mathbb{T}}(X_{1})\cong\mathrm{ker}(\delta_{0})=\langle\bar{\Gamma},\bar{M},X_{+}-X_{-}\rangle\cong\mathbb{Z}^{\oplus 3}$ (76)
$\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1})\cong\mathrm{coker}(\delta_{0})=\tfrac{\mathbb{Z}\langle\bar{e},\bar{h}\rangle}{\langle\bar{e}+\bar{h}\rangle}\cong\mathbb{Z}.$ (77)
@@
$0\to\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{2})\to\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1})\stackrel{{\scriptstyle\delta_{1}}}{{\to}}\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{2},X_{1})$ (78)
$\to\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{2})\to\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{1})$ ${\to}0,$ (79)
@@
$\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1},X_{0})\to\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(X_{1})$ $\stackrel{{\scriptstyle\delta_{1}}}{{\to}}\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{2},X_{1})$ (80)
@@
To determine $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{1})$ we see that this group is isomorphic to the kernel of the coboundary map:
$\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{0})\stackrel{{\scriptstyle}}{{\to}}$ $\mathcal{K}^{1}_{C_{4}\mathbb{T}}(X_{1},X_{0})$ (82)
@@
$\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{1})\cong\mathbb{Z}\oplus\mathbb{Z}$ (85)
@@
$\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}$ (86)
$\mathcal{K}^{-1}_{C_{4}\mathbb{T}}(T^{2})\cong 0.$ (87)
Therefore the bulk invariant of $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})$ is $\mathbb{Z}_{2}$, since it is the one coming from $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(X_{2},X_{1})$. To see that we can detect it with the Chern number of the spin-up bundle we just need to notice that we have the commutative square
@@
maps each component isomorphically to $\widetilde{K}^{0}(S^{2})$. Therefore the generator of the right hand side of Eqn. (92) can be detected with the generator of the spin-up component of the left hand side of Eqn. (92). Hence the bulk $\mathbb{Z}_{2}$ invariant of $\mathcal{K}^{0}_{C_{4}\mathbb{T}}(T^{2})$ is detected by the parity of the Chern number of the spin up bundle in $\mathcal{K}^{0}_{C_{4}\mathbb{T},S_{z}}(X_{2},X_{1}).$
@@
${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{2})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}$ (94)
@@
Summarizing: We have shown above in diagram (60) that the valence bundle $E$ of the Hamiltonian of Eqn. (1) has $1$ for spin-up Chern number. Therefore this bundle $E$ realizes the $\mathbb{Z}_{2}$ invariant of ${\mathcal{K}}^{0}_{C_{4}\mathbb{T}}(T^{2})$ and therefore it is topologically protected by this Chern number.

### v2

@@ III.2 $\mathbb{Z}_{2}$ bulk invariant for $C_{4}^{z}\mathbb{T}$ on 2D torus
$r\otimes\mathbb{Q}:\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})\otimes\mathbb{Q}\stackrel{{\scriptstyle\cong}}{{\to}}\left(K_{C_{2}^{z}}^{0}(T^{2})\otimes\mathbb{Q}\right)^{\mathbb{Z}_{2}}$ (46)
implies in particular that any integer invariant of $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$ could be detected with the integer invariant in $K_{C_{2}^{z}}^{0}(T^{2})^{\mathbb{Z}_{2}}$ given by the restriction map $r$.
The bundles that generate the bulk invariants of $K_{C_{2}^{z}}^{*}(T^{2})$ are well known and can be constructed. Let $\mathcal{H}$ be the line bundle over $T^{2}$ with $C_{2}^{z}$ action whose Chern number is $1$. We could take as $\mathcal{H}$ the line bundle given by the valence band of the Hamiltonian
@@
with $C_{2}^{z}$ action $(k_{x},k_{y})\mapsto(-k_{x},-k_{y})$ and associated matrix $\mathrm{diag}(i,-i)$. Here $C_{2}^{z}$ has for eigenvalue $i$ on $\Gamma$ and $-i$ on the other three TRIMS. The Chern number $c_{1}(\mathcal{H})$ of $\mathcal{H}$ is $1$. Its conjugate $\widehat{\mathcal{H}}$ has the opposite action of $C_{2}^{z}$, its Chern number $c_{1}(\widehat{\mathcal{H}})$ is $-1$ and the $C_{2}^{z}$ eigenvalues are the complex conjugate of the ones of $\mathcal{H}$.
The tensor power $\mathcal{H}^{\otimes 2n+1}$ of $\mathcal{H}$ is also a $C_{2}^{z}$-equivariant line bundle and the Chern number distinguishes the line bundles $\mathcal{H}^{\otimes 2n+1}$ in $K_{C_{2}^{z}}^{0}(T^{2})$.
@@
$\mathcal{H}^{\otimes 2n+1}\oplus\widehat{\mathcal{H}}^{\otimes 2n+1}\in K_{C_{2}^{z}}^{0}(T^{2})^{\mathbb{Z}_{2}}$ (48)
become all equal in $K_{C_{2}^{z}}^{0}(T^{2})$. Note that the Chern number of the sum $\mathcal{H}^{\otimes 2n+1}\oplus\widehat{\mathcal{H}}^{\otimes 2n+1}$ is zero, and they all share the same eigenvalues on the TRIMs. In other words, the bulk invariant of the bundles $\mathcal{H}^{\otimes 2n+1}\oplus\widehat{\mathcal{H}}^{\otimes 2n+1}$ is zero, and because of this, the Chern number for each separate piece cannot be extracted from the K-theory group.
@@
$\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2},$ (49)
where the $\mathbb{Z}_{2}$ invariant is precisely the bulk invariant that is of interest. This bulk invariant is precisely the Chern number (mod $2$) of one of the pieces of its decomposition in the generators of $K_{C_{2}^{z}}^{0}(T^{2})^{\mathbb{Z}_{2}}$. But since we cannot extract Chern numbers from $K_{C_{2}^{z}}^{0}(T^{2})^{\mathbb{Z}_{2}}$, we are going to add the symmetry of the spin $z$ operator in order to extract this Chern number.
Let $S_{z}$ be the spin $z$ and let us assume that the spin acts on the bundles defining $\mathcal{K}^{*}_{C_{4}^{z}\mathbb{T}}(T^{2})$. This happens for example when the Hamiltonian commutes with the spin. Incorporating $S_{z}$ to the group of symmetries we have that $\{S_{z},C_{4}^{z}\mathbb{T}\}=0$ and that $[S_{z},C_{2}^{z}]=0$. Now let us consider the following commutative diagram of restrictions:
where the left vertical maps forget the $S_{z}$ action, the left horizontal maps are the restriction to the complex part defined in Eqn. (39), and the right vertical map is the Chern number of the spin-up part. Here $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T},S_{z}}(T^{2})$ and $K^{0}_{C_{2}^{z},S_{z}}(T^{2})$ denote the equivariant magnetic and complex K-theories associated to the groups $\langle C_{4}^{z}\mathbb{T},S_{z}\rangle$ and $\langle C_{2}^{z},S_{z}\rangle$ respectively.
@@
Denote by $E$ the vector bundle that the valence bands of the altermagnetic Hamiltonian of Eqn. (1) defines. We know that for $G^{2}<M_{0}<2$ (say $M_{0}=1.5$), its restriction to the K-theory of $C_{2}^{z}$ splits as the line bundles defined by the valence bands for each Hamiltonian block; in other words
$r(E)=\mathcal{H}\oplus\widehat{\mathcal{H}}\in K_{C_{2}^{z}}^{0}(T^{2})^{\mathbb{Z}_{2}}.$ (55)
@@
Now we can finish our argument. In order to detect the Chern number of one of the pieces of the class $\mathcal{H}\oplus\widehat{\mathcal{H}}$ in the K-theory ${K}_{C_{2}^{z}}^{0}(T^{2})$ we use the spin $z$ splitting.
Remark: Note that we cannot extract the Chern number of one of the pieces in $\mathcal{H}\oplus\widehat{\mathcal{H}}$ from its image in ${K}_{C_{2}^{z}}^{0}(T^{2})$. In ${K}_{C_{2}^{z}}^{0}(T^{2})$ the element $\mathcal{H}\oplus\widehat{\mathcal{H}}$ has no bulk invariant.
Incorporating the spin, we see that the vector bundles in $K^{0}_{C_{2}^{z},S_{z}}(T^{2})$ can be split into spin-up and spin-down thus having:
$K^{0}_{C_{2}^{z},S_{z}}(T^{2})\cong K^{0}_{C_{2}^{z}}(T^{2})^{\uparrow}\oplus K^{0}_{C_{2}^{z}}(T^{2})^{\downarrow}.$ (61)
Since the $\mathbb{Z}_{2}$ action on vector bundles on $K^{0}_{C_{2}^{z},S_{z}}(T^{2})$ flips the spin (this follows from the fact that $C_{4}^{z}\mathbb{T}$ anticommutes with $S_{z}$), the $\mathbb{Z}_{2}$-invariant part can be parametrized with the spin-up bundle:
$K^{0}_{C_{2}^{z},S_{z}}(T^{2})^{\mathbb{Z}_{2}}\cong K^{0}_{C_{2}^{z}}(T^{2})^{\uparrow}.$ (62)
@@
This splitting allow us to extract only the component of the spin-up, and from this component we may take its Chern number; this is the spin-up Chern number. Knowing that this spin-up Chern number is odd, permits us to distinguish the bulk invariant of the bundle in $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$.
From diagrams (54) and (60) we can conclude the following. The bundle $E$ generates an integer invariant in $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T},S_{z}}(T^{2})$ since it maps to an integer invariant in ${K}^{0}_{C_{2}^{z},S_{z}}(T^{2})^{\mathbb{Z}_{2}}$. This follows from the fact that the restriction map in the upper left of diagram (54) induces an isomorphims rationally.
We now claim that the bundle $E$ generates the $\mathbb{Z}_{2}$-bulk invariant in $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$. For this we need to determine explicitly the complete K-theory invariants in ${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$. This will be done in the next section.
@@ III.2.1 Subdivision of the torus with 4 squares
To find the magnetic equivariant K-theory invariants in ${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$ we will make use of an equivariant subdivision of the torus. The long exact sequence of the pairs will allows to find the desired K-theory groups.
@@
$\mathcal{K}^{even}_{C_{4}^{z}\mathbb{T}}(X_{0})\cong$ $\mathbb{Z}\langle\bar{\Gamma},\bar{M},\bar{X}_{+},\bar{X}_{-}\rangle$ (66)
where both $\bar{\Gamma}$ and $\bar{M}$ denote the 2-dimensional representations of $\langle C_{4}^{z}\mathbb{T}\rangle$ on $\Gamma$ and $M$ respectively and $\bar{X}_{+}$ and $\bar{X}_{-}$ denote the 1-dimensional representations of $C_{2}^{z}$ in $X$ (recall that $(C_{2}^{z})^{2}=-1$).
@@
$\mathcal{K}^{odd}_{C_{4}^{z}\mathbb{T}}(X_{1},X_{0})\cong$ $\mathbb{Z}\langle\bar{e},\bar{h}\rangle$ (67)
$\mathcal{K}^{even}_{C_{4}^{z}\mathbb{T}}(X_{1},X_{0})\cong$ 00 0 (68)
$\mathcal{K}^{even}_{C_{4}^{z}\mathbb{T}}(X_{2},X_{1})\cong$ $\mathbb{Z}\langle\bar{Z}\rangle$ (69)
$\mathcal{K}^{odd}_{C_{4}^{z}\mathbb{T}}(X_{2},X_{1})\cong$ 00 0 (70)
which follows from the fact that the $C_{4}^{z}\mathbb{T}$ action on the 1 and 2-cells is free.
@@
$0\to\mathcal{K}^{-2}_{C_{4}^{z}\mathbb{T}}(X_{1})\to\mathcal{K}^{-2}_{C_{4}^{z}\mathbb{T}}(X_{0})\stackrel{{\scriptstyle\delta_{0}}}{{\to}}\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1},X_{0})$ (71)
$\to\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1})$ $\to 0.$
@@
$\mathcal{K}^{-2}_{C_{4}^{z}\mathbb{T}}(X_{0})\stackrel{{\scriptstyle\delta_{0}}}{{\to}}$ $\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1},X_{0})$ (72)
@@
$\mathcal{K}^{-2}_{C_{4}^{z}\mathbb{T}}(X_{0})=\widetilde{\mathcal{K}}_{C_{4}^{z}\mathbb{T}}(\Sigma^{2}X_{0})$ (75)
and therefore we can imagine the generators $\bar{\Gamma},\bar{M},\bar{X}_{+},\bar{X}_{-}$ to be virtual bundles over $S^{2}$. Since the conjugation action on $K^{0}(S^{2})$ maps the Hopf bundle to its inverse, we have that the restriction map $\mathcal{K}^{-2}_{C_{4}^{z}\mathbb{T}}(\{\Gamma,M\})\to{K}^{-2}(\{\Gamma,M\})$ is trivial. The image of $\bar{X}_{+}$ and $\bar{X}_{-}$ depends on the orientation of the 1-cells that attach to the point $X$. The 1-cell $h$ leaves from $X$ and the 1-cell $e^{\prime}$ arrives, but we need to use the automorphism that maps the information on $e^{\prime}$ to $e$ by conjugation, that on this dimension is given by multiplication by $-1$ (again the Hopf bundle is mapped to its inverse). Therefore we have
$\mathcal{K}^{-2}_{C_{4}^{z}\mathbb{T}}(X_{1})\cong\mathrm{ker}(\delta_{0})=\langle\bar{\Gamma},\bar{M},X_{+}-X_{-}\rangle\cong\mathbb{Z}^{\oplus 3}$ (76)
$\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1})\cong\mathrm{coker}(\delta_{0})=\tfrac{\mathbb{Z}\langle\bar{e},\bar{h}\rangle}{\langle\bar{e}+\bar{h}\rangle}\cong\mathbb{Z}.$ (77)
@@
$0\to\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{2})\to\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1})\stackrel{{\scriptstyle\delta_{1}}}{{\to}}\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{2},X_{1})$ (78)
$\to\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{2})\to\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{1})$ ${\to}0,$ (79)
@@
$\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1},X_{0})\to\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(X_{1})$ $\stackrel{{\scriptstyle\delta_{1}}}{{\to}}\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{2},X_{1})$ (80)
@@
To determine $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{1})$ we see that this group is isomorphic to the kernel of the coboundary map:
$\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{0})\stackrel{{\scriptstyle}}{{\to}}$ $\mathcal{K}^{1}_{C_{4}^{z}\mathbb{T}}(X_{1},X_{0})$ (82)
@@
$\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{1})\cong\mathbb{Z}\oplus\mathbb{Z}$ (85)
@@
$\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}$ (86)
$\mathcal{K}^{-1}_{C_{4}^{z}\mathbb{T}}(T^{2})\cong 0.$ (87)
Therefore the bulk invariant of $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$ is $\mathbb{Z}_{2}$, since it is the one coming from $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(X_{2},X_{1})$. To see that we can detect it with the Chern number of the spin-up bundle we just need to notice that we have the commutative square
@@
maps each component isomorphically to $\widetilde{K}^{0}(S^{2})$. Therefore the generator of the right hand side of Eqn. (92) can be detected with the generator of the spin-up component of the left hand side of Eqn. (92). Hence the bulk $\mathbb{Z}_{2}$ invariant of $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$ is detected by the parity of the Chern number of the spin up bundle in $\mathcal{K}^{0}_{C_{4}^{z}\mathbb{T},S_{z}}(X_{2},X_{1}).$
@@
${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})\cong\mathbb{Z}^{\oplus 2}\oplus\mathbb{Z}_{2}$ (94)
@@
Summarizing: We have shown above in diagram (60) that the valence bundle $E$ of the Hamiltonian of Eqn. (1) has $1$ for spin-up Chern number. Therefore this bundle $E$ realizes the $\mathbb{Z}_{2}$ invariant of ${\mathcal{K}}^{0}_{C_{4}^{z}\mathbb{T}}(T^{2})$ and therefore it is topologically protected by this Chern number.